OpenStudy (anonymous):

Find the mass and Volume of S2Cl2 at 1atm and 270K. If the total number of atoms are 1.879*10^26.

OpenStudy (anonymous):

@ParthKohli @JFraser Please help

OpenStudy (anonymous):

@countonme123 Someone please?

OpenStudy (anonymous):

Yes so i have 1.879*10^26 x 6.022*10^23 / 4 = 2.83 *10^49 mol This doesn't seem right.

OpenStudy (anonymous):

Then i got 312. 02 moles

OpenStudy (jfraser):

once you know how many total moles are present, use the molar mass to find out how much those moles \(weigh\) \[ mass = moles * (\frac{grams}{mole})\]

OpenStudy (photon336):

I agree with @jfraser and @rushwr you need to first find out how many moles you have of that sample. then finding the volume is easy just manipulating the formula. you know pressure you know temperature and number of moles. All you need is V